A faulty thermometer has LFP=10degree and UFP=110degree. What is actual temperature of an object if this thermometer reads 40degree celcius
 

Grade:11

2 Answers

Arun
4 years ago
 
Here faulty reading (x)=40
Lower fixed point of faulty readind (L.F.P)=10
Upper fixed point of faulty reading (U.F.P.)=110
Fundamental interval (F.I.)=110- 10 = 100
Correct reading(C)=?
c/100*(x-L.F.P)/F.I.
C/100=40/100 =0.4
C=0.4*100=40 celsius scale
Vikas TU
4 years ago
Dear student:
usually the thermometer fixed points basically are:
Lower point =0°c
Upper point=100°c
so in between there are 100 divisions.
each one division corresponds to 1°c

But in the above mentioned case :
Lower point=10°c
upper point=110°c
so total division=110-10=100
thus, 1 degree= 1 division=100/100
Now let us find the given temperature:
measured temperature 40 °c
=100x40/100
=40°c
Hope it is clear.